# SynCo: Synthetic Hard Negatives in Contrastive Learning for Better Unsupervised Visual Representations

### v3

@@ 5.1 Implementation Details
We pretrain SynCo on ImageNet ILSVRC-2012 [10] and its smaller ImageNet-100 subset [38] for ablation studies using a ResNet-50 [18] encoder. Our implementation builds upon MoCo-v2 [8]. For training, unless stated otherwise, we use $K=65$k ($K=16$k for ImageNet-100). For SynCo, we also have a warm-up of 10 epochs, i.e. for the first epochs we do not synthesize hard negatives. We set SynCo’s hyperparameters $\sigma$, $\delta$, and $\eta$ to 0.01. For hard negative generation, we use the top $N=1024$ hardest negatives, with $N_{1}=N_{2}=N_{3}=256$ and $N_{4}=N_{5}=N_{6}=64$. For ImageNet linear evaluation, we train a linear classifier on frozen features for 100 epochs, using a batch size of 256 and a cosine learning rate schedule. Initial learning rates are set to 30.0 for ImageNet and 10.0 for ImageNet-100. To evaluate transfer learning, we apply SynCo to object detection tasks. For PASCAL VOC [12], we fine-tune a Faster R-CNN [35] (R50-C4) on trainval07+12 and test on test2007. For COCO [28], we use a Mask R-CNN [19] (R50-C4), fine-tuning on train2017 and evaluating on val2017. We employ Detectron2 [46] and report standard AP metrics, following [20] without additional hyperparameter tuning. Detailed implementation details are provided in the supplementary material.
@@ 5.2 Linear Evaluation on ImageNet
PIRL 63.6 -
LA 60.2 -
InfoMin 70.1 89.4
SimSiam 68.1 -
SimCLR 69.3 -
MoCo 60.7 -
MoCo-v2 67.5 90.1
MoCHI 66.9 -
@@
PIRL 800 63.6 -
InfoMin 800 73.0 91.1
BYOL 1000 74.3 91.6
SimSiam 800 68.1 -
SimCLR 1000 69.3 -
Barlow Twins 1000 73.2 91.0
MoCo-v2 800 71.1 90.1
MoCHI 800 68.7 -
@@ 5.3 Semi-supervised Training on ImageNet
InstDis 200 - - 39.2 77.4
PIRL 800 30.7 60.4 57.2 83.8
SimCLR 1000 48.3 65.6 75.5 87.8
Barlow Twins 1000 55.0 69.7 79.2 89.3
BYOL 1000 53.2 68.8 78.4 89.0
SwAV 800 53.9 70.2 78.5 89.9
@@ 5.4 Transferring to Detection
InstDis - - - 37.4 57.6 40.6 34.1 54.6 36.4
PIRL 55.5 81.0 61.3 38.5 57.6 41.2 34.0 54.6 36.2
InfoMin 57.6 82.7 64.6 39.0 58.5 42.0 34.1 55.2 36.3
SwAV 56.1 82.6 62.7 38.2 58.2 41.3 33.8 55.2 35.9
SimSiam 57.0 82.4 63.7 39.2 59.3 42.1 34.4 56.0 36.7
Barlow Twins 56.8 82.6 63.4 39.2 59.0 42.5 34.3 56.0 36.5
MoCo 55.5 81.5 61.3 38.5 58.3 41.6 33.6 54.8 35.6
MoCo-v2 57.0 82.4 63.6 39.0 58.6 41.9 34.2 55.4 36.2
MoCHI 57.5 82.7 64.4 39.2 58.9 42.4 34.4 55.5 36.6
@@ 5.5 Ablation Study
Type of hard negative. We evaluate the impact of each synthetic hard negative type on pretraining. For this, we select the top $N=1024$ hardest negatives and generate $N_{i}=256$, $i=1,2,\ldots,6$ negatives. We train SynCo without hard negatives (equivalent to MoCo-v2) for 100 epochs and measure top-1 and top-5 accuracy. Subsequently, we train SynCo using each type of hard negative individually, and then using all six types in combination. The results of these ablations are presented in Figure 6(a). We see that every SynCo configuration outperform the MoCo-v2 baseline.
Parameter sensitivity analysis. We conducted ablations on the parameters $\sigma$, $\delta$, and $\eta$ of SynCo’s type 4, type 5, and type 6 negatives, respectively. The results, presented in Figure 6(b), show that varying these parameters does not lead to significant differences in performance. This suggests that SynCo is robust across a wide range of values for $\sigma$, $\delta$, $\eta$.
Queue size. We investigate the effect of queue size $\mathcal{Q}$ on performance. We train SynCo and MoCo-v2 with reduced queue sizes. Our results, presented in Figure 6(c), reveal that SynCo performs comparably to MoCo-v2 across various queue sizes. With smaller queues, SynCo underperforms compared to MoCo-v2. This can be attributed to the fact that the total generated synthetic negatives are too hard for the task and harm performance, a finding that is also observed in [23]. However, as the queue increases, SynCo performs on par with MoCo-v2. At the largest queue size tested, SynCo outperforms MoCo-v2.
@@ 6 Conclusion
Broader Impact
The presented research should be categorized as research in the field of unsupervised learning. This work may inspire new algorithms, theoretical, and experimental investigation. The algorithm presented here can be used for many different vision applications and a particular use may have both positive or negative impacts, which is known as the dual use problem. Besides, as vision datasets could be biased, the representation learned by SynCo could be susceptible to replicate these biases.

### v4

@@ 5.1 Implementation Details
We pretrain SynCo on ImageNet ILSVRC-2012 [11] and its smaller ImageNet-100 subset [39] for ablation studies using a ResNet-50 [19] encoder. Our implementation builds upon MoCo-v2 [9]. For training, unless stated otherwise, we use $K=65$k ($K=16$k for ImageNet-100). For SynCo, we also have a warm-up of 10 epochs, i.e. for the first epochs we do not synthesize hard negatives. We set SynCo’s hyperparameters $\sigma$, $\delta$, and $\eta$ to 0.01. For hard negative generation, we use the top $N=1024$ hardest negatives, with $N_{1}=N_{2}=N_{3}=256$ and $N_{4}=N_{5}=N_{6}=64$. For ImageNet linear evaluation, we train a linear classifier on frozen features for 100 epochs, using a batch size of 256 and a cosine learning rate schedule. Initial learning rates are set to 30.0 for ImageNet and 10.0 for ImageNet-100. To evaluate transfer learning, we apply SynCo to object detection tasks. For PASCAL VOC [13], we fine-tune a Faster R-CNN [36] on trainval07+12 and test on test2007. For COCO [29], we use a Mask R-CNN [20], fine-tuning on train2017 and evaluating on val2017. We employ Detectron2 [47] and report standard AP metrics, following [21] without additional hyperparameter tuning. Detailed implementation details are provided in the supplementary material.
@@ 5.2 Linear Evaluation on ImageNet
PIRL [31] 63.6 -
LA [56] 60.2 -
InfoMin [40] 70.1 89.4
SimSiam [8] 68.1 -
SimCLR [7] 69.3 -
MoCo [21] 60.7 -
MoCo-v2 [9] 67.5 90.1
MoCHI [24] 66.9 -
@@
PIRL [31] 800 63.6 -
InfoMin [40] 800 73.0 91.1
BYOL [15] 1000 74.3 91.6
SimSiam [8] 800 68.1 -
SimCLR [7] 1000 69.3 -
BT [52] 1000 73.2 91.0
MoCo-v2 [9] 800 71.1 90.1
MoCHI [24] 800 68.7 -
@@ 5.3 Semi-supervised Training on ImageNet
InstDis [48] 200 - - 39.2 77.4
PIRL [31] 800 30.7 60.4 57.2 83.8
SimCLR [7] 1000 48.3 65.6 75.5 87.8
BT [52] 1000 55.0 69.7 79.2 89.3
BYOL [15] 1000 53.2 68.8 78.4 89.0
SwAV [5] 800 53.9 70.2 78.5 89.9
@@ 5.4 Transferring to Detection
InstDis [48] - - - 37.4 57.6 40.6 34.1 54.6 36.4
PIRL [31] 55.5 81.0 61.3 38.5 57.6 41.2 34.0 54.6 36.2
InfoMin [40] 57.6 82.7 64.6 39.0 58.5 42.0 34.1 55.2 36.3
SwAV [5] 56.1 82.6 62.7 38.2 58.2 41.3 33.8 55.2 35.9
SimSiam [8] 57.0 82.4 63.7 39.2 59.3 42.1 34.4 56.0 36.7
BT [52] 56.8 82.6 63.4 39.2 59.0 42.5 34.3 56.0 36.5
MoCo [21] 55.5 81.5 61.3 38.5 58.3 41.6 33.6 54.8 35.6
MoCo-v2 [9] 57.0 82.4 63.6 39.0 58.6 41.9 34.2 55.4 36.2
MoCHI [24] 57.5 82.7 64.4 39.2 58.9 42.4 34.4 55.5 36.6
@@ 5.5 Ablation Study
Ablation on type of hard negative.
We evaluate the impact of each synthetic hard negative type on pretraining. For this, we select the top $N=1024$ hardest negatives and generate $N_{i}=256$, $i=1,2,\ldots,6$ negatives. We train SynCo without hard negatives (equivalent to MoCo-v2) for 100 epochs and measure top-1 and top-5 accuracy. Subsequently, we train SynCo using each type of hard negative individually, and then using all six types in combination. The results of these ablations are presented in Figure 5(a). We see that every SynCo configuration outperform the MoCo-v2 baseline.
Ablation on hyperparameters.
We conducted ablations on the parameters $\sigma$, $\delta$, and $\eta$ of SynCo’s type 4, type 5, and type 6 negatives, respectively. The results, presented in Figure 5(b), show that varying these parameters does not lead to significant differences in performance. This suggests that SynCo is robust across a wide range of values for $\sigma$, $\delta$, $\eta$.
Ablation on queue size.
We investigate the effect of queue size $\mathcal{Q}$ on performance. We train SynCo and MoCo-v2 with reduced queue sizes. Our results, presented in Figure 5(c), reveal that SynCo performs comparably to MoCo-v2 across various queue sizes. With smaller queues, SynCo underperforms compared to MoCo-v2. This can be attributed to the fact that the total generated synthetic negatives are too hard for the task and harm performance, a finding that is also observed in [24]. However, as the queue increases, SynCo performs on par with MoCo-v2. At the largest queue size tested, SynCo outperforms MoCo-v2.
@@ 6 Conclusion
Acknowledgments
We would like to express our gratitude to Andreas Floros for his valuable feedback, particularly his assistance with equations, notations, and insightful discussions that greatly contributed to this work.